rssLink RSS for all categories
 
icon_orange
icon_green
icon_red
icon_orange
icon_red
icon_green
icon_green
icon_orange
icon_green
icon_blue
icon_green
icon_green
icon_green
icon_red
icon_orange
icon_red
icon_green
icon_green
icon_red
icon_red
icon_green
icon_green
icon_red
icon_blue
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_green
icon_red
 

FS#31796 — gra1-sd5a-n9

Attached to Project— Reseau Internet et Baies
Incident
Gravelines   → GRA-1
CLOSED
100%
Le routeur gra1-sd5a-n9 ne répond plus.
Nous investiguons.

---

The gra1-sd5a-n9 router no longer responds.
We're investigating.
Date:  Thursday, 24 May 2018, 19:15PM
Reason for closing:  Done
Additional comments about closing:  le routeur a été mis à jour et le trafic est de nouveau dessus. Tout est ok maintenant

------

Router is up-to-date and all the traffic is going throught it. Everything is fine now
Comment by OVH - Thursday, 24 May 2018, 16:31PM

Tous les vlans ont été basculés sur le router b.
Notre fournisseur est en train d'analyser la situation.

---

All vlans have been moved to router b.
Our supplier is analyzing the situation.


Comment by OVH - Thursday, 24 May 2018, 17:30PM

Nous venons de reloader le routeur A.
Nous préparons l'upgrade du routeur A.

---

We just reloaded router A.
We are preparing the upgrade of router A.


Comment by OVH - Thursday, 24 May 2018, 17:54PM

Nous mettons à jour le routeur côté A, tout le trafic passant par le côté B il n'y a donc pas d'impact
----
We are upgrading the router on A side. All the traffic is going through the A side so there is no impact


Comment by OVH - Thursday, 24 May 2018, 18:16PM

Le routeur A est maintenant à jour, nous allons remettre le trafic dessus
---
Router A is up-to-date we will put back the traffic on it


Comment by OVH - Thursday, 24 May 2018, 19:05PM

Le trafic est de nouveau sur le routeur A, il y a des perturbations sur 2000 IP le temps de la bascule (environ 10 min). Tout est OK maintenant
------
Traffic is back on the A side, there were some perturbation on 2000 IP during the swap (10 min). Everything is OK now